1. The market demand function for a product is given by Q=300-2P. How much consumer surplus do they

receive when

P=45?

P=30?

2. The market supply curve for a product is given by P=10+2Qs. How much producer surplus do they

receive when

P=18

P=30

3. Let the market supply and demand curves are given by the equations

P= 40 + 4Qs and P= 100 – 2Qd. What is the consumer and producer surplus at the equilibrium price and quantity?

4. Suppose the market demand and supply functions are given by Qd=60-P and Qs=P-20. Determine the consumer and producer surplus if a price ceiling of 32 is imposed in this market. What is the amount of dead weight loss?

5. The domestic demand and supply and demand curves

for petrol are as follows: Supply: P=50+ Demand: P=200-20 What will be the consumer and producer surplus if a price ceiling of 75 is imposed?
1
Expert's answer
2020-12-17T07:39:11-0500

1. Q = 300 - 2P, then P = 150 - 0.5Q.

If P = 45, then "CS = 0.5*(150 - 45)*210 = 11,025."

If P = 30, then "CS = 0.5*(150 - 30)*240 = 14,400."

2. P = 10 + 2Qs, then Q = 0.5P - 5.

If P = 18, then "PS = 0.5*(18 - 10)*4 = 16."

If P = 30, then "PS = 0.5*(30 - 10)*10 = 100."

3. P = 40 + 4Qs and P = 100 – 2Qd.

In equilibrium Qd = Qs and Pd = Ps, so:

40 + 4Q = 100 – 2Q,

Q = 10 units,

P = 100 - 2*10 = 80.

"CS = 0.5*(100 - 80)*10 = 100."

"PS = 0.5*(80 - 40)*10 = 200."

4. If a price ceiling of 32 is imposed in this market, then both the consumer and producer surplus will decrease and the dead weight loss will occur.

5. If a price ceiling of 75 is imposed, then both the consumer and producer surplus will decrease and the dead weight loss will occur.